# Color: A Framework for Applying Graph Coloring to Subgraph Cardinality Estimation

### v1

@@ 8.3. Inference Latency
Fig. 4 shows the distribution of inference latencies for each method across workloads. We can see that the inference latency of COLOR lies in the middle of the competing methods across workloads. On the smaller, less cyclic queries of human, it achieves a very fast median latency of around $10^{-4}$ seconds due to partial aggregation, and on the larger more complex queries of patents it has a median latency of .$\sim.05$ seconds via sampling.
Further, when compared to the other graph summarization methods, SumRDF and BSK++, the methods from our framework scale far better to larger queries. The former methods timeout (¿1 minute) on queries of even moderate size as they consider the exponential number of potential colorings of the query. This occurs even when using partial aggregation (as in BSK++), demonstrating the necessity of sampling to achieve consistent latencies.
@@ 8.4. Statistics Size & Build Time
Graph summarization approaches allow for a smooth tradeoff between accuracy and size/build time; a more granular summary of the graph will take more space but more accurately capture the structure of the data graph. Fortunately, even a compact summary ($<20$MB) can be highly accurate as shown by Fig. 7. Part of this compactness comes from the fact that we store our summary sparsely. This means that if two colors do not share an edge with a particular attribute then we don’t explicitly store any degree statistic about this combination. Due to this, a better coloring can actually result in a more compact summary because many of these combinations won’t occur if the nodes have been properly partitioned. On Human and Lubm80, AlleyTPI’s pattern index requires 600 and 300 MB, respectively.
@@ 8.5. Updates
In Fig. 11, we evaluate the effectiveness of our method for handling updates (Sec. 7.3). To do this, we randomly partition the edges of the human dataset into an initial data graph and an ensuing set of updates, and we construct a lifted graph using the former then update it by adding one edge/vertex at a time based on the latter. Intuitively, when more of the graph is provided at the beginning, the lifted graph will be more accurate because it can take advantage of that knowledge when coloring the graph. Additionally, due to uncertainty about the attributes for added edges and vertices, the proposed update method generalizes and may update statistics where it may be unnecessary, leading to errors at high update proportions.
We find that, as more of the lifted graph is updated, the accuracy remains very consistent and degrades to the traditional independence estimator. This implies that a full rebuilding of the lifted graph can occur very infrequently and be amortized over many updates.
Due to the small size of the lifted graph, updating it is very fast. The latency for vertex updates was roughly 0.4 ms while for edge updates it was roughly 0.1 ms.
Figure 11. Relative Error vs Proportion Updated (Human)
@@ 8.6. Micro-Benchmarks
Fig. 13 demonstrates the effects of partial aggregation and sampling. In this figure, using the Youtube data graph, we observe how inference latency changes across different query pathwidths if partial aggregation or sampling is used. Recall that pathwidth is a measure of cyclicity where a query with pathwidth 1 is acyclic and higher pathwidth queries are increasingly interconnected. Without partial aggregation, the estimation begins to time out (¿1 min) when query pathwidths exceed 2. This is because higher pathwidth queries tend to be larger and this naive approach has an exponential runtime w.r.t. the size of the query. When using partial aggregation without sampling, the estimation times out after query pathwidths exceed 4. Lastly, when sampling is applied, the inference latency becomes linear in the size of the query and unrelated to the pathwidth. The results demonstrate the importance of including partial aggregation for achieving speedups without affecting accuracy, and that the use of sampling can achieve consistent, fast inference.
@@
In Fig. 14, we vary the sample size used during inference on the Youtube workload to demonstrate the efficacy of our sampling method. Using a very small sample size results in significant underestimation, but even a moderate number of samples quickly converges to the accuracy of a large number of samples. Further, this figure shows that importance sampling speeds up this convergence significantly over a naive uniform sample. For example, the performance at $250$ samples for importance sampling is roughly equal to the accuracy of uniform sampling at $1000$ samples.
@@
Figure 13. Build Time vs Query Path Width (Youtube)
@@ 9. Conclusion
We develop COLOR, a framework for producing lifted graph summaries from colorings. We define inference over the lifted graph for acyclic and cyclic queries, developing optimizations to accelerate estimation. We empirically validate COLOR’s superior performance on eight benchmark datasets, comparing it to state-of-the-art methods. COLOR is up to $10^{3}\times$ more accurate than the baselines and stands out for never experiencing estimation failure. It gracefully handles updates, degrading to only 3×3$\times$3 × worse error when half of the data is replaced.
@@ Acknowledgements.
Appendix A Proof of Thm. 1
In this section, we prove the following theorem for simple graphs, but the extension to property graphs is straightforward.
Theorem 3.
(9) $|\hom(Q,G)|=\Phi(Q,\mathcal{G})$
Proof.
We start by noting that each match in the data graph, $\pi\in Hom(Q,G)$, is associated with precisely one coloring, $\pi^{\prime}\in Hom(Q,F)$, based on the matched vertices’ colors, i.e. $\pi^{\prime}=\sigma_{S}\circ\pi$. We denote the set of matches with this coloring as $Hom(Q,G|\pi^{\prime})$. If we calculate $|Hom(Q,G|\pi^{\prime})|$ for each coloring, $\pi^{\prime}$, then we can compute the total as $\sum_{\pi^{\prime}\in Hom(Q,F)}|Hom(Q,G|\pi^{\prime})|$. By this logic and equation 4, we simply need to show that $W(\pi^{\prime})=|Hom(Q,G|\pi^{\prime})|$.
We note that $Q$ is a acyclic and proceed inductively on the topological ordering $v_{1},\ldots,v_{|Q|}$. Let $Q_{i}$ be the sub-tree restricted to the vertices $v_{1},\ldots,v_{i}$, and let $\pi^{\prime}_{i}$ be the coloring restricted to these vertices. We begin with the base case,
$W(\pi^{\prime}_{1})=\psi(\pi^{\prime}_{1}(v_{1}))=|Hom(Q_{1},G|\pi^{\prime}_{1})|$
The number of matches to a vertex of a color is equal to the number of vertices in that color, so this is immediately true. Now, suppose that $W(\pi^{\prime}_{i})=|Hom(Q_{i},G|\pi^{\prime}_{i})|$, and we will show that $W(\pi^{\prime}_{i+1})=|Hom(Q_{i+1},G|\pi^{\prime}_{i+1})|$. We restate Eq. 4 as follows,
$W(\pi^{\prime}_{i+1})=\psi(\pi^{\prime}_{i+1}(v_{1}))\prod_{(v_{j},v_{k})\in E_{Q_{i+1}}}\tau((\pi^{\prime}_{i+1}(v_{j}),\pi^{\prime}_{i+1}(v_{k})))$
Let $v_{j}$ be the parent of $v_{i+1}$, and we can express $W_{\psi,\tau}(\pi^{\prime}_{i+1})$ inductively,
$W(\pi^{\prime}_{i+1})=W(\pi^{\prime}_{i})\cdot\tau((\pi_{i+1}(v_{j}),\pi_{i+1}(v_{i+1})))$
By our inductive assumption, this means,
$W(\pi^{\prime}_{i+1})=|Hom(Q_{i},G|\pi^{\prime}_{i})|\cdot\tau((\pi_{i+1}(v_{j}),\pi_{i+1}(v_{i+1})))$
Denote the number of edges that a vertex $v\in V_{G}$ has to vertices with color $C$ as $deg_{G}(v|C)$. Because $\sigma_{S}$ is a stable coloring, we know,
$deg_{G}(v|C)=deg_{G}(v^{\prime}|C)\,\,\,\forall v,v^{\prime}\in V_{G}\,\,s.t.\,\,\sigma_{S}(v)=\sigma_{S}(v^{\prime})$
By the definition of $\tau$, this degree is precisely,
$deg_{G}(v|C)=\tau(\sigma_{S}(v_{j}),C)$
We denote this degree with its color as $deg_{G}(C|C^{\prime})$. Returning to our expression for $W_{\psi,\tau}(\pi^{\prime}_{i+1})$, we can now plug in our degree expression,
$W(\pi^{\prime}_{i+1})$ $=|Hom(Q_{i},G|\pi^{\prime}_{i})|\cdot deg_{G}(\pi^{\prime}_{i+1}(v_{j})|\pi^{\prime}_{i+1}(v_{i+1}))$
By simply expanding $|Hom(Q_{i},G|\pi^{\prime}_{i})|$ into a sum, we get,
$W^{std}(\pi^{\prime}_{i+1})$ $=\sum_{\pi_{i}\in Hom(Q_{i},G|\pi^{\prime}_{i})}deg_{G}(\pi^{\prime}_{i+1}(v_{j})|\pi^{\prime}_{i+1}(v_{i+1}))$
At this point, we use the fact that $deg(\pi^{\prime}_{i}(v_{j})|\pi^{\prime}_{i+1}(v_{i+1}))$
$=deg(\pi_{i}(v_{j})|\pi^{\prime}_{i+1}(v_{i+1}))$ when $\sigma_{S}(\pi_{i}(v_{j}))=\pi^{\prime}_{i}(v_{j})$, which is assured by $\pi\in Hom(Q_{i},G|\pi^{\prime}_{i})$.
$W^{std}(\pi^{\prime}_{i+1})$ $=\sum_{\pi_{i}\in Hom(Q_{i},G|\pi^{\prime}_{i})}deg_{G}(\pi_{i}(v_{j})|\pi^{\prime}_{i+1}(v_{i+1}))$
Because $Q$ is a tree, the RHS is the target of the induction,
$W^{std}(\pi^{\prime}_{i+1})=|Hom(Q_{i+1},G|\pi^{\prime}_{i+1})|$
Appendix B Proof of Theorem 2
We begin by restating the theorem,
Theorem 4.
Given a query graph, $Q$, a lifted graph, $F$, and a decomposable estimator $W$, $\Phi(Q,F,W_{\psi,\tau})$ can be computed in time $O(|C|^{tw(Q)})$ where $tw(Q)$ is the treewidth of $Q$ where $\Phi(Q,F,W_{\psi,\tau})$ is defined,
$\Phi(Q,F,W_{\psi,\tau})=\sum_{c_{v_{1}},...,c_{v_{|Q|}}\in C}\psi(c_{0})\prod_{i=1}^{|E_{Q}|}\omega(\pi_{v_{k}\rightarrow c_{v_{k}}},e_{i}|e_{1},...,e_{i-1})$
We now define tree decompositions,
Definition 0.
Given a graph $H$, a tree decomposition $T(E_{T},V_{T},\chi,\gamma)$ is composed of four pieces,
  1. (1)
    $E_{T}$, $V_{T}$ are the edges and vertices of a tree
  2. (2)
    $\chi:V_{T}\rightarrow 2^{V_{H}}$ is a function which maps vertices in the tree to sets of vertices in $H$
  3. (3)
    $\gamma:V_{T}\rightarrow 2^{E_{H}}$ is a function which maps vertices in the tree to sets of edges in $H$
Lastly, it has two requirements,
  1. (1)
    For all $v\in V_{H}$, the set of vertices in $V_{T}$ that contain $v$ form a connected sub-tree
  2. (2)
    Each edge in $E_{H}$ is mapped to exactly one vertex of $V_{T}$ by $\gamma$ and the vertex of $V_{T}$ which it is mapped to includes both the endpoints
The treewidth is then defined as follows,
Definition 0.
Given a graph $H$, let the set of valid tree decompositions be $\mathcal{T}_{H}$. The treewidth is then defined as,
$\min_{T\in\mathcal{T}_{H}}\max_{v\in V_{T}}|\chi(v)|-1$
Intuitively, graphs that are ”more acyclic” will have a lower treewidth and graphs that are ”more cyclic” have a higher treewidth.
Proof.
For this problem, we can take advantage of tree decompositions by using them to structure the summation in (8). Suppose that the treewidth of $Q$ is $k$ and let $T(E_{T},V_{T},\chi)$ be a tree decomposition which matches this width. To avoid confusion, we will denote vertices of $T$ as $v_{i}^{\prime}$ and vertices of $Q$ as $v_{i}$. Further, let $v_{1}^{\prime},\cdots,v_{|V_{T}|}^{\prime}$ be a topological ordering of $T$ where $v_{1}^{\prime}$ is the root of the tree, and let $v_{1},\cdots,v_{|V_{Q}|}$ be an ordering of $Q$ such that the sub-tree corresponding to $v_{i}$ is not a sub-tree of $v_{>i}$. Let $Par(v^{\prime})$ denote the parent of $v^{\prime}$ in $T$ and let $Ch(v^{\prime})$ be the set of child vertices of $v^{\prime}$ in $T$. Lastly, we denote the set of query vertices which associated with $v_{|V_{T}|}^{\prime}$ and not its parent as $\mathbf{X}_{i}=\chi(v_{i}^{\prime})\setminus\chi(Par(v_{i}^{\prime}))$. We denote the set of query vertices which are associated with $v_{|V_{T}|}^{\prime}$ AND its parent as $\mathbf{Y}_{|V_{T}|}=\chi(v_{|V_{T}|}^{\prime})\cap\chi(Par(v_{|V_{T}|}^{\prime}))$.
We will proceed iteratively on these vertices starting with $v_{|V_{T}|}^{\prime}$ and proceeding backwards towards the root.
Base Case: Our base case is $v_{|V_{T}|}^{\prime}$ which is necessarily a leaf of the true due to the topological ordering. We denote the output of the base case as the function $S_{|V_{T}|}(c_{y_{1}},\ldots,c_{y_{m}})$ and define it as follows,
$S_{|V_{T}|}(c_{y_{1}},\ldots,c_{y_{m}})=\sum_{c_{x_{1}},\ldots,c_{x_{n}}\in C}\prod_{e_{i}\in\gamma(v_{|V_{T}|}^{\prime})}\omega(\pi_{x_{j}\rightarrow c_{x_{j}}},e_{i}|e_{1},\ldots,e_{i-1})$
At this point of the computation, we will fully materialize the values of the function $S_{|V_{T}|}$ which has a domain of size $|C|^{|Y_{|V_{T}|}|}$ and each point requires computing a summation of $|C|^{|X_{|V_{T}|}|}$ terms. Therefore, the entire computation requires $|C|^{|\chi(v_{|V_{T}|}^{\prime})|}\leq|C|^{k}$ time.
Inductive Case: Suppose that all vertices $v_{j}^{\prime}$ where $j>i$ have already been processed, we now consider handling $v_{i}^{\prime}$. The output of this step is defined as,
$S_{i}(c_{Y_{v_{i}^{\prime},1}},\ldots,c_{Y_{v_{i}^{\prime},m}})=\sum_{c_{X_{v_{i}^{\prime},1}},\ldots,c_{X_{v_{i}^{\prime},n}}\in C}$ $\prod_{e_{i}\in\gamma(v_{i}^{\prime})}\omega(\pi_{X_{v_{i}^{\prime},j}\rightarrow c_{X_{v_{i}^{\prime},j}}},e_{i}|e_{1},\ldots,e_{i-1})\cdot$
$\prod_{v_{j}^{\prime}\in Ch(v_{i}^{\prime})}S_{v_{j}^{\prime}}(c_{Y_{v_{j}^{\prime},1}},\ldots,c_{Y_{v_{j}^{\prime},m}})$
As in the base case, the computation required to fully materialize the values of $S_{i}$ is bounded by $|C|^{|\chi(v_{|V_{T}|}^{\prime})|}\leq|C|^{k}$.
At this point, we just need to confirm that $S_{1}$ is equal to $\Phi(Q,F,W_{\psi,\tau})$ by showing that the sequence of computations represents a valid rearrangement of the sums of the original formula. To this end, we note that every query vertex appears in the summation of precisely one inductive step. This is due to the fact that each query vertex forms a sub-tree in the tree decomposition, so the root of that sub-tree is the only tree vertex which contains it and whose parent does not. Further, each instance of $\omega(\pi_{X_{v_{i}^{\prime},j}\rightarrow c_{X_{v_{i}^{\prime},j}}},e_{i}|e_{1},\ldots,e_{i-1})$ occurs precisely once because of the requirement that each edge of the query graph occurs in one vertex of the tree decomposition.

### v2

@@ 8.3. Inference Latency
Fig. 4 shows the distribution of inference latencies for each method across workloads. We can see that the inference latency of COLOR lies in the middle of the competing methods across workloads. On the smaller, less cyclic queries of human, it achieves a median latency of around $10^{-4}$ seconds due to partial aggregation, and on the more complex queries of patents it has a median latency of .$\sim.05$ seconds via sampling. Similar to prior work, we select a one minute timeout because cardinality estimators may be called many times during the query planning phase (Wang et al., 2021; Kim et al., 2021; Zhao et al., 2023; Park et al., 2020b).
Compared to other graph summarization methods, SumRDF and BSK++, our framework scales far better to larger queries. The former methods timeout on queries of even moderate size as they consider the exponential number of potential colorings of the query. This occurs even when using partial aggregation (as in BSK++), demonstrating the necessity of sampling to achieve consistent latencies.
@@ 8.4. Statistics Size & Build Time
Graph summarization approaches allow for a smooth tradeoff between accuracy and size/build time; a more granular summary of the graph will take more space but more accurately capture the structure of the data graph. Even a compact summary ($<20$MB) can be highly accurate as shown by Fig. 7. This comes from the fact that we store our summary sparsely. If two colors do not share an edge with a particular attribute, then we don’t explicitly store any degree statistic about this combination. A better coloring can actually result in a more compact summary because many of these combinations won’t occur if the nodes have been properly partitioned. On Human and Lubm80, AlleyTPI’s pattern index requires 600 and 300 MB, respectively. LSS takes a considerable time to train across all datasets. This is the average training time over all folds of the cross validation, and, crucially, it does not include the time necessary to collect the training data (i.e. to run the queries and calculate a true cardinality).
@@ 8.5. Updates
In Fig. 11, we evaluate the effectiveness of our method for handling updates (Sec. 7.3). To do this, we randomly partition the edges of the human dataset into an initial data graph and an ensuing set of updates, and we construct a lifted graph using the former then update it by adding one edge/vertex at a time based on the latter. Intuitively, when more of the graph is provided at the beginning, the lifted graph will be more accurate because it can take advantage of that knowledge when coloring the graph. We find that, as more of the lifted graph is updated, the accuracy remains very consistent and degrades to the traditional independence estimator. So, a full rebuilding of the lifted graph can occur very infrequently and be amortized over many updates. Due to the small size of the lifted graph, each update is very fast. The latency for vertex updates was roughly 0.4 ms while for edge updates it was roughly 0.1 ms.
@@ 8.6. Micro-Benchmarks
Fig. 13 demonstrates the effects of partial agg. and sampling. Using the Youtube dataset, we study the effect of partial agg. and sampling on inference latency across queries with different pathwidths. Recall that pathwidth is a measure of cyclicity where pathwidth 1 is acyclic and higher pathwidth queries are denser. Without partial agg., estimation times out (¿1 min) when pathwidth exceeds 2. Higher pathwidth queries are larger and the naive approach is exponential w.r.t. the query size. Using only partial agg., estimation times out when pathwidths exceed 4. Lastly, when sampling is applied, the inference latency becomes linear in the size of the query and unrelated to the pathwidth. The results demonstrate that partial agg. achieves speedups without affecting accuracy and sampling can achieve consistent, fast inference.
Figure 11. Relative Error vs Proportion Updated (Human)
@@
In Fig. 14, we vary sample size to demonstrate the efficacy of our sampling method. Using a very small sample results in significant underestimation, but even a moderate number of samples quickly converges to the accuracy of a large number of samples. Further, this figure shows that importance sampling speeds up this convergence significantly over a naive uniform sample. For example, the performance at $250$ samples for importance sampling is roughly equal to the accuracy of uniform sampling at $1000$ samples.
@@
Figure 13. Inference Time vs Query Pathwidth (Youtube)
@@ 9. Conclusion
We develop COLOR, a framework for producing lifted graph summaries from colorings. We define inference over the lifted graph for acyclic and cyclic queries, developing optimizations to accelerate estimation. We empirically validate COLOR’s superior performance on eight benchmarks and compare to state-of-the-art methods. COLOR is up to $10^{3}\times$ more accurate than the baselines and stands out for never experiencing estimation failure. It gracefully handles updates, degrading to only 3×3$\times$3 × worse error when half of the data is replaced.